In an achievement test, the mean score of normally distributed values is 70 and the standard deviation is 12. What is the percentage of students who got a score of 85 and above?







1
Expert's answer
2022-03-05T05:16:23-0500
"P(X\ge85)=1-P(X<85)"

"=1-P(Z<\dfrac{85-70}{12})=1-P(Z<1.25)"


"\approx0.10565"

"10.565\%"